OpenStudy (anonymous):

Solve by completing the square: p^2-3p-88=0

OpenStudy (jdoe0001):

do you know what a "perfect square trinomial" is?

OpenStudy (anonymous):

This is what I did so far: p^2-3p-88=0 +88 +88 p^2-3p=88 then (b/2)^2=(-3/2)^2 Then I got lost

OpenStudy (anonymous):

And i remember hearing it, but I forgot what it is

OpenStudy (jdoe0001):

hmmm does \(\large \bf (a\pm b)^2 = a^2\pm 2ab+b^2\) ring a bell?

OpenStudy (anonymous):

ummm wow, no not at all :/

OpenStudy (jdoe0001):

heheh... well... that's a "perfect square trinomial" notice the pattern 1st element squared middle is "2" times both 2nd element squared

OpenStudy (anonymous):

Right, but it says on this equation that I have to solve by completing the square, and that formulas not in the book

OpenStudy (jdoe0001):

well..... keep that in mind for one because "completing the square" more or less means just that

OpenStudy (anonymous):

I'm kinda confued

OpenStudy (jdoe0001):

\( \bf p^2-3p-88\implies p^2-3p=88\implies (p^2-3p+{\color{red}{ \square ?}}^2)=88\) any ideas what we need there to get a "perfect square trinomial"

OpenStudy (anonymous):

I was going to say 88, but I realized you moved that to the other side. Sorry I'm not sure

OpenStudy (anonymous):

Would it be 2?

OpenStudy (jdoe0001):

well..... notice in the parentheses we have 2 terms the 1st one and the middle one notice the middle one recall that the middle is 2 times both so...... what do you think would be the missing term, the 2nd one?

OpenStudy (anonymous):

You said the middle is 2 times both, so like 3*3 so 9? Sorry if that's wrong

OpenStudy (jdoe0001):

well... let's see the middle term is 3p <--- we know that much so we can say that, the 1st term is "p" so 2 * p * "2nd term" = 3p if you solve that for "2nd term".... you'd see what that is :)

OpenStudy (anonymous):

How do I solve for second term? I know p=1

OpenStudy (anonymous):

Would it be 6?

OpenStudy (jdoe0001):

\(\large \bf 2\cdot p\cdot \square =3p\implies \square =?\)

OpenStudy (anonymous):

I am not getting any of this. Are you sure this is completing the square, cause I didn't learn this in class?

OpenStudy (jdoe0001):

well...... when they mean "completing the SQUARE" is referring to a "perfect SQUARE trinomial" as shown above

OpenStudy (anonymous):

I don't know how to do that tho. Is 88 suppose to be in the blank?

OpenStudy (jdoe0001):

hmmm do you know how to do linear simplification? that is, solving for a term or a variable?

OpenStudy (anonymous):

umm I don't think so. I'm going to put a example up of how they should me how to do it, so hold on

OpenStudy (anonymous):

x^2+6x-3=0 +3 +3 x^2+6=3 (b/2)^2= (6/3)^2= 3^2= 9 Add to both sides x^2+6x+9=3+9 (x+3)^2=12 sqrt(x+3)^2=sqrt(12) x=3(+-)3.46 x=0.46 or x=-6.46

OpenStudy (anonymous):

If you can't help it's fine I give up lol

OpenStudy (jdoe0001):

hmmm say for if you had .... 2x + 3 = 5 could you solve that for "x"?

OpenStudy (anonymous):

2x+3=5 -3 -3 2x=2 x=1

OpenStudy (jdoe0001):

well... that's simplifying by linear simplification

OpenStudy (anonymous):

Right...But that's a little simpler then this because you only have one variable not 2

OpenStudy (jdoe0001):

the example you showed... above looks a bit convoluted let us find the middle term gimme a sec

OpenStudy (jdoe0001):

\(\bf p^2-3p-88\implies p^2-3p=88\implies (p^2-3p+{\color{red}{ \square ?}}^2)=88 \\ \quad \\ -------------------------------\\ {\color{purple}{ 2\cdot p\cdot \square =3p\implies 2p\square =3p\implies \square =\cfrac{3\cancel{ p }}{2\cancel{ p }}}}\qquad thus\\ -------------------------------\\ \left[p^2-3p+{\color{red}{ \left(\frac{3}{2}\right)}}^2\right]=88\implies \left(p-\frac{3}{2}\right)^2=88\qquad thus \\ \quad \\ p-\frac{3}{2}=\pm\sqrt{88}\implies p=\pm\sqrt{88}+\frac{3}{2}\)

OpenStudy (jdoe0001):

the 2nd term is 3/2 because if we make the 1st term "p" 2nd term "3/2" then the middle term is \(\bf 2\cdot \textit{1st term}\cdot \textit{2nd term}\implies \cancel{ 2 }\cdot p\cdot \frac{3}{\cancel{ 2 }}\implies 3p\)

OpenStudy (anonymous):

OMG! I actually get what you did there. Thanks for typing it out like that, it makes more scenes. So is that all you need or is there still more?

OpenStudy (jdoe0001):

that's.... all....once the perfect square is completed just take square root to both sides and add the 3/2

OpenStudy (jdoe0001):

hmmmm actiually wait a sec

OpenStudy (anonymous):

this is so hard :/ If you can't get it tho don't worry about it. This is the only one on the whole page I didn't get. I appreciate the help

OpenStudy (jdoe0001):

lemme correct myself some when we are completing the square we ADDED \(\left(\frac{3}{2}\right)^2\) but all we're really doing is "borrowing from our good friend Mr Zero" , 0 so if we add \(\left(\frac{3}{2}\right)^2\) we also have to SUBTRACT \(\left(\frac{3}{2}\right)^2\)

OpenStudy (jdoe0001):

so.... lemme redo that quick

OpenStudy (jdoe0001):

\(\bf p^2-3p-88\implies p^2-3p=88\implies (p^2-3p+{\color{red}{ \square ?}}^2)=88 \\ \quad \\ -------------------------------\\ {\color{purple}{ 2\cdot p\cdot \square =3p\implies 2p\square =3p\implies \square =\cfrac{3\cancel{ p }}{2\cancel{ p }}}}\qquad thus\\ -------------------------------\\ \left[p^2-3p+{\color{red}{ \left(\frac{3}{2}\right)}}^2\right]-{\color{red}{ \left(\frac{3}{2}\right)}}^2=88\implies \left(p-\frac{3}{2}\right)^2=88+{\color{red}{ \left(\frac{3}{2}\right)}}^2\qquad thus \\ \quad \\ p-\frac{3}{2}=\pm\sqrt{88+\frac{3}{2}}\implies p=\pm\sqrt{88+\frac{3}{2}}+\frac{3}{2}\)

OpenStudy (jdoe0001):

keep in mind, that we're only "borrowing" from Zero, is all we're doing so whatever amount we ADD, we also have to SUBTRACT it

OpenStudy (jdoe0001):

hmmm I missed the squared part... shoot lemme fix that

OpenStudy (jdoe0001):

\(\bf p^2-3p-88\implies p^2-3p=88\implies (p^2-3p+{\color{red}{ \square ?}}^2)=88 \\ \quad \\ -------------------------------\\ {\color{purple}{ 2\cdot p\cdot \square =3p\implies 2p\square =3p\implies \square =\cfrac{3\cancel{ p }}{2\cancel{ p }}}}\qquad thus\\ -------------------------------\\ \left[p^2-3p+{\color{red}{ \left(\frac{3}{2}\right)}}^2\right]-{\color{red}{ \left(\frac{3}{2}\right)}}^2=88\implies \left(p-\frac{3}{2}\right)^2=88+{\color{red}{ \left(\frac{3}{2}\right)}}^2\qquad thus \\ \quad \\ p-\frac{3}{2}=\pm\sqrt{88+\left(\frac{3}{2}\right)^2}\implies p=\pm\sqrt{88+\left(\frac{3}{2}\right)}+\frac{3}{2}\)

OpenStudy (jdoe0001):

hmmm \(\bf p^2-3p-88\implies p^2-3p=88\implies (p^2-3p+{\color{red}{ \square ?}}^2)=88 \\ \quad \\ -------------------------------\\ {\color{purple}{ 2\cdot p\cdot \square =3p\implies 2p\square =3p\implies \square =\cfrac{3\cancel{ p }}{2\cancel{ p }}}}\qquad thus\\ -------------------------------\\ \left[p^2-3p+{\color{red}{ \left(\frac{3}{2}\right)}}^2\right]-{\color{red}{ \left(\frac{3}{2}\right)}}^2=88\implies \left(p-\frac{3}{2}\right)^2=88+{\color{red}{ \left(\frac{3}{2}\right)}}^2\qquad thus \\ \quad \\ p-\frac{3}{2}=\pm\sqrt{88+\left(\frac{3}{2}\right)^2}\implies p=\pm\sqrt{88+\left(\frac{3}{2}\right)^2}+\frac{3}{2}\)

OpenStudy (jdoe0001):

anyhow... had the squred missing from the fraction in the radical

OpenStudy (anonymous):

Wow. Thanks again, you helped so much. I'm defiantly going to ask my teacher to go over this tomorrow.

OpenStudy (jdoe0001):

so... simplifying that we could say that

OpenStudy (jdoe0001):

\(\bf p=\pm\sqrt{88+\left(\frac{3}{2}\right)}+\frac{3}{2}\implies p=\pm \sqrt{88+\frac{9}{4}}+\frac{3}{2} \\ \quad \\ p=\sqrt{\cfrac{361}{4}}+\cfrac{3}{2}\implies p=\pm \cfrac{\sqrt{361}}{\sqrt{4}}+\cfrac{3}{2}\implies p=\pm \cfrac{19}{2}+\cfrac{3}{2} \\ \quad \\ p=\pm \cfrac{\cancel{ 22 }}{\cancel{ 2 }}\to \pm 11\)

OpenStudy (jdoe0001):

darn... it.. missed the square again anyhow \(\bf p=\pm\sqrt{88+\left(\frac{3}{2}\right)^2}+\frac{3}{2}\implies p=\pm \sqrt{88+\frac{9}{4}}+\frac{3}{2} \\ \quad \\ p=\sqrt{\cfrac{361}{4}}+\cfrac{3}{2}\implies p=\pm \cfrac{\sqrt{361}}{\sqrt{4}}+\cfrac{3}{2}\implies p=\pm \cfrac{19}{2}+\cfrac{3}{2} \\ \quad \\ p=\pm \cfrac{\cancel{ 22 }}{\cancel{ 2 }}\to \pm 11\)